Circular No. 5815

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

CH CYGNI A. Skopal and M. F. Bode, Liverpool John Moores University; A.

Evans, University of Keele; and Z. Urban, D. Chochol, and L. Hric, Astronomical Institute, Skalnate Pleso Observatory, communicate: “There has been an increase in the activity of the symbiotic binary CH Cyg (cf. IAUC 5504). Low-dispersion SWP and LWP, and high-dispersion LWP, spectra were obtained with the International Ultraviolet Explorer from the VILSPA station of the European Space Agency on 1992 Nov. 24 and 1993 May 28. The level of the continuum increased approximately by a factor of two, from 1.2 to 2.5 × 10^{−12} erg cm^{−2} s^{−1} A^{−1} around 150 nm (and at the limit of the spectrum at 330 nm), and from 0.4 to 0.9 × 10^{−12} erg cm^{−2} s^{−1} A^{−1} at around 240 nm, before any reddening correction was applied. The line spectrum was characterized by absorption features, which were strong on the later SWP spectrum, and emissions of Lyα, O IV] + Si IV (140 nm), C IV (155 nm), O III] (166 nm), Si III] (189 nm), C III] (191 nm), C II] (233 nm), and [O II] (247 nm). The high-resolution spectra exhibited numerous Fe II lines in absorption. The Mg II 279.6- and 280.3-nm doublet emission lines are each cut by two strong absorption components. Photoelectric photometry carried out at Skalnate Pleso confirmed the recent increase in activity, particularly in the U band: 1992 Nov. 18.96 UT, U = 9.496, B = 9.638, V = 8.242; Dec. 23.75, 8.184, 8.944, 8.014; 1993 Apr. 17.99, 7.55, 8.49, 7.82; June 1.90, 7.57, 8.39, 7.69.”

V1974 CYGNI G. S. Rossano, R. J. Rudy, R. C. Puetter, and D. K. Lynch, Space

and Environment Technology Center, The Aerospace Corporation, write: “We have observed this object on June 3.5 UT with the Lick Observatory 3-m Shane telescope and the Aerospace cooled scanning germanium spectrometer. A spectrogram (range 750-1350 nm, resolution 300-550) shows strong He I (1083.0 nm), weak Paβ through Paϵ, and weak lines of He II, [O II] (733.1 nm), and [N I] (1040.0 nm). Also seen are the coronal lines [Fe XI] at 789.2 nm, [S VIII] at 991.1 nm, and [S IX] at 1252.0 nm. [S IX] is now stronger than Paβ. [Fe XIII] at 1074.7 and 1079.8 nm appear blended with the He I 1083.0-nm feature. Unidentified features at 1110, 1190, and 1200 nm observed in 1992 August and October are still present. A monochromatic continuum magnitude of 12.8 at 1250 nm was observed.”

1993 June 15 Daniel W. E. Green